OpenStudy (anonymous):

Find the indicated nth term of the geometric sequence. 6th term: a[5]=3/256, a[8]=-3/16,384

OpenStudy (anonymous):

Does a[5] mean that it is the 5th term?

OpenStudy (anonymous):

what is 384?

OpenStudy (anonymous):

Assuming that a[5] means the 5th term, I am going to make it the 1st term and a[8] the 4th term. going by the Geometric Sequence formula a[n]=a1]*r^(n-1) , where r is the common ratio, then a[4]=a[1]*r*(4-1) -3/16/384=3/256*r^3 Multiplying both sides by 256/3, the result is -1/64=r^3 Now finding the cubic root on both side , the result is r=-1/4. to find the sixth term (or in this case the second term) using the formula again a[6] equivalent to a[2]=a[1]*(-1/4)^(2-1) a[6] =(3/256)*(-1/4)=-3/1024

OpenStudy (anonymous):

oh ok thank you! this makes perfect sense!